A 0.25 kg billiard ball traveling at 1 m/s along the positive x axis strikes an identical stationary billiard ball at the origin. After the collision the original ball has a velocity making an angle of 25 degrees with the x axis and the stationary ball has a velocity making an angle of -10 degrees with the x axis. What is the x component of the total momentum after the collision? O 0.5 kg m/s 0.25 kg m/s O 0 kg m/s 1 kg m/s
Answer #1

Given

m1 = m2 = m = 0.25 kg

u1 = 1 m/s

u2 = 0

Now using conservation of momentum,

Total x-component of momentum before collision = Total x-component of momentum after collision

Total x-component of momentum after collision = m1*u1 + m2*u2 = 0.25*1 + 0.25*0 = 0.25 kg m/s